F(x)=x^2+pc+q a,b are roots then a^2 – b^2
______________________________________________________________________________________________________________________________________________

Grade:10

1 Answers

manoj kumar
6 years ago
For finding roots F(x)=0x^2+pc+q=0Root. a= +(pc+q)^1/2 b= -(pc+q)^1/2 Then a^2 = pc+q b^2 = pc+q Then a^2- b^2 = pc+q-pc-q = 0